# ORCID: 0009-0007-8992-9682

### v2

@@ 3.3.3 Perturbation on Syntax
We use six perturbations from ReCode (Wang et al., 2022) where three of them: DeadCode Insertion, Operand Swap, and For-While-Switch are generated by NatGen (Chakraborty et al., 2022). Additionally the rest three of them: Variable Renaming by CodeBERT (Feng et al., 2020), Variable Renaming with NatGen style, and generate random variable names using a combination of alphabetic and numeric characters are prepared by ReCode itself.
The following list explains the syntax perturbations used in this work. Table 4 illustrates examples of how the perturbations affect the original code snippet. The changes that happen in each of the syntax perturbations are highlighted.
@@ 3.3.4 Perturbation on Format
By now, we have proposed a total of 28 strategies for perturbation. And in our experiments, for each coding problem, we apply each of the 28 strategies to the original prompt and generate five perturbed prompts. Therefore, for each coding problem, in total, we have 28*5 perturbed prompts, and for each perturbed prompt, one code is generated.
@@ 3.4.1 Robust Pass ($RP_{s}@k$)
Pass@k is a commonly used metric for code generation tasks. As we are generating multiple perturbed prompts from a single prompt, we use a modified version of this metric. From the initial prompt, there is $s$ number of perturbed prompts. Additionally, $k$ generations are made from each perturbed prompt, for a total of $s*k$ generations from a single prompt. In this work, we use $s=5$ and $k=1$. We use the $RP_{s}@k$ metrics as equation 2.
@@ 3.4.2 Robust Drop ($RD_{s}@k)$
To compare the robustness of the generated code from perturbed data with the generated code from nominal data, the Robust Drop metric is used. This metric states the fall in the performance compared to the nominal data. In equation 3, this metric is defined.
@@ 4.1 Dataset
As stated in Section 3.2, we utilized two datasets to evaluate the robustness of code generation models. Below, we further detail the datasets used in our experiments.
HumanEval-X. HumanEval-X [^4] https://huggingface.co/datasets/THUDM/humaneval-x is an extended version of the HumanEval [^5] https://github.com/openai/human-eval dataset, originally designed to evaluate code generation models in Python. HumanEval-X expands this benchmark to support multiple programming languages, including C++, Java, JavaScript, and Go. It consists of 164 prompts, each accompanied by a canonical solution and language-specific test cases. In our study, we leverage HumanEval-X to assess code translation and bug classification across Java, C++, and JavaScript. The dataset provides a total of X test cases, with an average of Y test cases per language, ensuring comprehensive evaluation across different programming paradigms.
EvalPlus. As the second dataset, we use EvalPlus after the curation process described in Section 3.2. The prompts remain consistent with those in HumanEval-X, except for cases where we made adjustments to accommodate corner cases. Additionally, EvalPlus provides more extensive and critical test cases. However, many HumanEval-X solutions were not directly suitable for EvalPlus test cases. To address this, we manually wrote canonical solutions for each programming language in HumanEval-X and then tested them using the EvalPlus-converted test cases. This adaptation ensures that the evaluations are meaningful across multiple programming languages. The statistics of the test cases are shown in Table 3
@@ 4.2 Studied Code Language Model
In our experiment, we focus on evaluating four distinct code generation models. Specifically, we select Incoder-1b, Incoder-6b, CodeGen-2b-Multi, and CodeGen-6b-Multi as they have been trained in multiple programming languages. These models are specifically chosen due to their varying number of parameters, allowing us to assess our experiments across a range of model complexities. Due to the excessive size and our inability to run them on our machine, we exclude certain large models, such as CodeGen-16B-Multi, from our experiment.
@@ 4.3 Setup
[…]
@@ 4.4.1 Sentence Similarity Measurement
For the Java, C++, and JavaScript datasets, we assess the semantic similarity between the perturbed and non-perturbed prompts. To do this, we employ the sentence cosine similarity measurement, following the same methodology as ReCode. We calculate the similarities between the perturbed and nominal datasets using the sentence embeddings from MPNet (Song et al., 2020). Specifically, we utilize MPNet embeddings to measure the cosine similarities for the docString and function name sections. For docstrings, we directly compare them after extraction, while for function names, we split them into lists of words and then compare them. Table 4 shows the summary of similarities, and Table 5 shows the similarities for individual perturbations.
@@
Categories Perturbations Java JavaScript CPP docString ChangeCharCase 1.0 1.0 1.0 TenseTransformationPast 0.98 0.98 0.98 EnglishInflectionalVariation 0.95 0.95 0.95 WhitespacePerturbation 0.89 0.9 0.89 ButterFingersPerturbation 0.86 0.86 0.87 TenseTransformationFuture 0.97 0.97 0.97 SynonymSubstitution 0.88 0.88 0.88 SwapCharactersPerturbation 0.9 0.89 0.9 SynonymInsertion 0.9 0.89 0.91 BackTranslation 0.94 0.93 0.92 Function FuncRenameSynonymSub 0.72 0.75 0.72 FuncRenameButterFinger 0.56 0.62 0.57 FuncRenameSwapChar 0.83 0.83 0.82 FuncRenameSnake/CamelCase 0.89 0.89 0.89 FuncRenameInflectionalVariation 0.93 0.94 0.93 FuncRenameChangeChar 0.37 0.41 0.36
@@ 4.4.2 Code Similarity Measurement
To assess the similarities between the perturbed and nominal data, we utilize CodeBLEU’s Syntax and DataFlow scores (Ren et al., 2020). For our analysis, we adapt the CodeBLEU score calculation method employed by CodeXGlue (Lu et al., 2021). Additionally, to calculate scores for C++, we incorporate NaturalCC (Wan et al., 2022) since the DataFlowGraph for C++ is not available in CodeXGlue, which is essential for computing the DataFlow score. The summary of similarity scores is presented in Table 4, while Table 6 displays the individual similarity scores. In general, the scores are higher, indicating significant similarity, with the exception of DeadCodeInserter. The inclusion of a code chunk within the code block leads to the dissimilarity between the perturbed prompt and the nominal prompt.
@@ 4.4.3 Human Annotation of Naturalness and Semantic Similarity in Perturbed Prompts
We conduct a random sampling of 300 perturbed and non-perturbed pairs, comprising 100 Java, 100 C++, and 100 JavaScript prompts. Each prompt is annotated by a single annotator, who provides three values: the naturalness of the perturbed prompt, the naturalness of the non-perturbed prompt, and the semantic similarity between the perturbed and non-perturbed prompts. The naturalness score ranges continuously from 0 to 1, where 0 indicates unnatural text and 1 signifies highly natural text. Similarly, the semantic similarity score also ranges from 0 to 1, where 0 represents a complete change in meaning, 0.5 indicates partial changes, and 1 denotes exact semantic preservation. We adopted the annotation methodology employed by ReCode (Wang et al., 2022), which utilized five annotators. However, for our study, we used one annotator to annotate the 300 samples. The findings of the annotations are presented in Table 7. Table 7 summarizes the human annotation results across Java, C++, and JavaScript, and also includes a comparison with ReCode’s Python-based annotations. The results show that non-perturbed prompts maintain near-perfect naturalness scores (above 0.98), while perturbed prompts still retain relatively high naturalness (around 0.8–0.85). Semantic similarity scores are also consistently high (above 0.94), indicating that the perturbations generally preserve meaning. These findings suggest that although the perturbations slightly reduce naturalness, they largely maintain the original semantics, validating the use of these perturbations in robustness evaluations.
Table 7: Result of Human Annotations for 300 Random Samples in Three Different Languages
Scope Java CPP JavaScript ReCode (Python)
Naturalness (Nominal) $\uparrow$ 0.99 0.98 0.98 0.92
Naturalness (Perturbed) $\uparrow$ 0.85 0.81 0.8 0.75
Semantic Similarity $\uparrow$ 0.94 0.96 0.95 0.92
@@ 5 Experiment Results
In this section, we aim to answer three research questions (RQs). For each RQ, we present the motivation and the approach to addressing the RQ, and the corresponding results.
RQ1: How robust are large language models (LLMs) across different programming languages?
@@
Motivation: Our experiment encompasses three distinct languages and four different models. A prior investigation (Wang et al., 2022) demonstrates the influence of perturbations on the robustness of code generation models. Nonetheless, the question of whether language-specific variations impact robustness remains unanswered. Hence, in this study, we aim to address this gap by exploring discrepancies in Large Language Models’ (LLMs) robustness across diverse programming languages (Java, C++, and JavaScript), shedding light on the extent of disparities among these languages.
@@
Approach: To address this research question, we perform experiments using Java, C++, and JavaScript across four models: Incoder-1B, Incoder-6B, CodeGen-Multi-2B, and CodeGen-Multi-6B. We evaluate the performance using three metrics: $RP_{s}@k$ measures robustness, $RD_{s}@k$ assesses the decrease in robustness due to changes, and $RR_{s}@k$ captures the relative differences between the original and modified datasets. The perturbations related to the Function Name and the DocString are compared with the nominal prompts (prompts that have only DocString and method signature), while the perturbations related to Syntax and Format are compared with the partial prompts (prompts that have some lines of code along with DocString and method signature)
@@
Results: Overall, the performance of models deteriorates after applying the perturbations. Table 8 presents the worst-case performance of four models—Incoder-1B, Incoder-6B, CodeGen-2B-Multi, and CodeGen-6B-Multi—under different perturbation types (DocString, Function Signature, Syntax, and Format) across three programming languages (Java, C++, and JavaScript). Similarly, Table 9 shows the results in partial-case scenarios, where a model passes at least 3 out of 5 generated samples.
@@
Across both tables, we observe a consistent performance drop after applying perturbations. For instance, in Table 8, the metric $\text{RD}_{5}@1$, which quantifies the worst-case degradation, shows severe drops—up to 0.9 (90%) in some DocString perturbation cases for Incoder-1B (Java). Even more superficial perturbations, such as Syntax and Format, lead to noticeable performance degradation, albeit to a lesser extent. Table 9 highlights that even under relaxed partial-pass criteria, the relative performance drop (RR5@1) remains substantial, typically ranging from 5% to 20%, depending on the model and language. This suggests that perturbations—whether semantic or syntactic—can significantly impact model behavior and reliability.
@@
Interestingly, we observe a slight performance improvement in the partial-case scenario for Java under DocString and Function perturbations. This suggests that in some cases, modifying the function or the NL description can yield better results, hinting at the potential of prompt refinement as a strategy for improving code generation performance.
@@
In general, all models demonstrate performance degradation under perturbations, reaffirming the brittleness of current LLM-based code generation systems. While models such as CodeGen-6B show slightly better robustness, the consistent drop across all settings underscores the need for robustness-aware evaluation and training. Moreover, the variation in impact across programming languages suggests that the models’ sensitivity may be language-dependent, warranting further investigation into language-specific robustness strategies.
@@
If we see the language-wise trends, the three languages—Java, C++, and JavaScript—exhibit distinct behaviors under input perturbations. Java shows the highest robustness, with notable RP5@1 gains and minimal RD5@1 drops, particularly under Format and Syntax perturbations. Its structured syntax likely helps models infer intent despite surface-level changes.

### v4

@@ 3.3.3 Perturbation on Syntax
We use the same six perturbations used by ReCode (Wang et al., 2022) where three of them: DeadCode Insertion, Operand Swap, and For-While-Switch are generated by NatGen (Chakraborty et al., 2022). The rest three of them: Variable Renaming by CodeBERT (Feng et al., 2020), Variable Renaming with NatGen style, and generate random variable names using a combination of alphabetic and numeric characters are prepared by ReCode itself.
The following list explains the syntax perturbations used in this work. Figure 5 illustrates examples of how the perturbations affect the original code snippet. The changes that happen in each of the syntax perturbations are highlighted.
@@ 3.3.4 Perturbation on Format
We applied 29 perturbation strategies in total: 7 perturbations on function name, 10 on docstring, 6 syntax perturbation, and 6 formatting perturbations. For each coding problem, 28 of these strategies were used to alter the original prompt. The remaining strategy depends on the language (camelCase is used for C++/JavaScript, whereas snake_case is used for Java). Each strategy produces five perturbed versions of the prompt, resulting in 28 × 5 perturbed prompts per problem, and we generate one code sample for each perturbed prompt. Some strategies may occasionally generate identical text, but using a non-zero sampling temperature during generation still produces varied model outputs across runs.
We follow ReCode’s practice of generating five perturbed prompts per strategy to capture the natural variability introduced by stochastic perturbation methods. Multiple samples provide a more stable and representative estimate of each strategy’s robustness, as some techniques exhibit high variability while others are largely deterministic.
For certain strategies, especially when perturbing very short text (e.g., single-word function names), some or all of the five perturbed prompts may become identical. Deterministic methods such as back-translation can even produce identical outputs across all five samples. We treat such duplication as an inherent property of the strategy, reflecting its limited perturbation diversity rather than a methodological issue.
Identical perturbed prompts do not affect our robustness results; they reflect the limited variation that some perturbation methods naturally produce.
@@ 3.4.1 Robust Pass ($RP_{s}@k$)
Pass@k is a commonly used metric for code generation tasks. As we are generating multiple perturbed prompts from a single prompt, we use a modified version of this metric. From the initial prompt, there is $s$ number of perturbed prompts. Additionally, $k$ generations are made from each perturbed prompt, for a total of sks*k generations from a single prompt. In this work, we use $s=5$ and $k=1$. We use the $RP_{s}@k$ metrics as equation 2.
@@
We follow recent evaluation practices that report pass@1 as the primary metric for code generation (Chen et al., 2021; Nijkamp et al., 2022), and therefore use $k=1$ for each prompt. Our goal is to evaluate the model’s robustness to input variation rather than its sampling variance. Accordingly, instead of generating multiple outputs from a single prompt, we generate five perturbed versions of each sample. Multiple generations from the same prompt would primarily capture decoding randomness, whereas multiple perturbed prompts reveal how sensitive the model is to small input changes. Using five perturbed prompts thus provides a more direct and meaningful robustness assessment.
@@ 3.4.2 Robust Drop ($RD_{s}@k)$
The Robust Drop metric is used to measure the decrease in robustness of the generated code under perturbed data compared to the original data. This metric states the fall in the performance compared to the nominal data. In equation 3, this metric is defined.
@@ 4.1 Dataset
As described in Section 3.2, we construct a new dataset by combining, extending, and converting two existing datasets to evaluate the robustness of code generation models. Below, we provide a detailed description of these source datasets and our final curated dataset used in the experiments.
HumanEval-X. HumanEval-X [^4] https://huggingface.co/datasets/THUDM/humaneval-x is an extended version of the HumanEval [^5] https://github.com/openai/human-eval dataset, originally designed to evaluate code generation models in Python. HumanEval-X expands this benchmark to support multiple programming languages, including C++, Java, JavaScript, and Go. It consists of 164 prompts, each accompanied by a canonical solution and language-specific test cases. In our study, we leverage HumanEval-X to assess code translation and bug classification across Java, C++, and JavaScript. This dataset provides a total of 820 test cases for 820 problems, with an average of 164 test cases per programming language and approximately 7.7 unit tests per problem.
EvalPlus. EvalPlus Liu et al. (2023) is an enhanced version of HumanEval, originally developed for Python. It introduces a substantially larger set of corner cases and refined prompts that are better aligned with the associated test cases. In this work, we adapt the EvalPlus test cases for our three target programming languages, following the curation process described in Section 3.2. The prompts are largely consistent with those in HumanEval-X, except in instances where minor modifications were required to accommodate language-specific corner cases.
While EvalPlus offers more comprehensive and rigorous test coverage, many HumanEval-X reference solutions were not directly compatible with its test cases. To address this, we manually implemented canonical solutions for each target language and evaluated them using the converted EvalPlus test cases. This adaptation ensures that our cross-language evaluations are both valid and comparable. The detailed statistics of the resulting test cases are presented in Table 3.
@@ 4.2 Studied Code Language Model
In our experiment, we focus on evaluating six distinct code generation models: Incoder-1b, Incoder-6b, CodeGen-2b-Multi, and CodeGen-6b-Multi Magicoder-7B, and QwenCode-2.5-6.7B as they have been trained in multiple programming languages. These models are specifically chosen due to their varying number of parameters, allowing us to assess our experiments across a range of model complexities. Due to the excessive size and our inability to run them on our machine, we exclude certain large models, such as CodeGen-16B-Multi, StarCoder-13B, etc, from our experiment.
@@ 4.3 Setup
Model Setup and Hyperparameter Selection. We use HuggingFace’s AutoModelForCausalLM API to load pretrained code generation models for all target programming languages. We did not perform any training or fine-tuning; we directly evaluated them in their pretrained form. Our goal is to assess the performance of existing models under consistent experimental conditions.
The same set of hyperparameters is applied across models, chosen through a small grid search on a held-out subset of the evaluation data. The temperature is set to 0.2 to reduce randomness and increase the likelihood of generating syntactically valid and test-passing code. The nucleus sampling parameter (top_p) is set to 0.95 to retain diversity while suppressing low-probability outputs. These values are taken from the ReCode Wang et al. (2022). The maximum generation length is set to 1536 tokens to prevent truncation. This value is based on our experience that it provides sufficient output while keeping inference time low.
We follow the same practice of pre-processing, prompting, and post-processing steps from prior works (Nijkamp et al., 2022) (Wei et al., 2023) (Hui et al., 2024). Language and model-specific stopping rules are applied to keep generations consistent and clean. Extra content, such as redundant helper functions or repeated class definitions, is removed. For Java, generation stops at the appearance of a second Solution class. For JavaScript and C++, it stops at the second docstring or comment block.
@@ 4.4.1 Sentence Similarity Measurement
For the Java, C++, and JavaScript datasets, we assess the semantic similarity between the perturbed and non-perturbed prompts. To do this, we employ the sentence cosine similarity measurement, following the same methodology as ReCode. We calculate the similarities between the perturbed and nominal datasets using the sentence embeddings from MPNet (Song et al., 2020). To be more specific, we utilize MPNet embeddings to measure the cosine similarities for the docString and function name sections. For docstrings, we directly compare them after extraction, while for function names, we split them into lists of words and then compare them. Table 4 shows the summary of similarities, and Table 5 shows the similarities for individual perturbations. These scores range from 0 to 1, where values close to 0 indicate dissimilarity and values close to 1 indicate high similarity. The two tables show that the scores are generally close to 1, indicating that the perturbations only slightly modify the docstrings and function names while largely preserving their content.
@@
Categories Perturbations Java JavaScript CPP docString ChangeCharCase 1.0 1.0 1.0 TenseTransformationPast 0.98 0.98 0.98 EnglishInflectionalVariation 0.95 0.95 0.95 WhitespacePerturbation 0.89 0.9 0.89 ButterFingersPerturbation 0.86 0.86 0.87 TenseTransformationFuture 0.97 0.97 0.97 SynonymSubstitution 0.88 0.88 0.88 SwapCharactersPerturbation 0.9 0.89 0.9 SynonymInsertion 0.9 0.89 0.91 BackTranslation 0.94 0.93 0.92 Function FuncRenameSynonymSub 0.72 0.75 0.72 FuncRenameButterFinger 0.56 0.62 0.57 FuncRenameSwapChar 0.83 0.83 0.82 FuncRenameSnake/CamelCase 0.89 0.89 0.89 FuncRenameInflectionalVariation 0.93 0.94 0.93 FuncRenameChangeChar 0.87 0.91 0.86
@@ 4.4.2 Code Similarity Measurement
To assess the similarities between the perturbed and nominal data, we utilize CodeBLEU’s Syntax and DataFlow scores (Ren et al., 2020). For our analysis, we adapt the CodeBLEU score calculation method employed by CodeXGlue (Lu et al., 2021). Additionally, to calculate scores for C++, we incorporate NaturalCC (Wan et al., 2022) since the DataFlowGraph for C++ is not available in CodeXGlue, which is essential for computing the DataFlow score. The summary of similarity scores is presented in Table 4, while Table 6 displays the individual similarity scores. These scores range from 0 to 1 as well, where values closer to 1 indicate higher similarity between the perturbed and nominal code. Overall, the scores are relatively high across most perturbations, suggesting that the code semantics and structure are largely preserved. However, the DeadCodeInserter perturbation yields comparatively lower scores, as the insertion of additional code blocks introduces noticeable structural differences that reduce similarity.
@@ 4.4.3 Human Annotation of Naturalness and Semantic Similarity in Perturbed Prompts
We conducted a random sampling of 300 prompt pairs, comprising 100 Java, 100 C++, and 100 JavaScript examples, each containing a perturbed and non-perturbed version of the same prompt. Each pair was annotated independently by two annotators, who provided three scores:
  • the naturalness of the perturbed prompt
  • the naturalness of the non-perturbed prompt
  • the semantic similarity between the perturbed and non-perturbed versions.
We follow 5-point Likert scale during the annotation process, both annotators provide discrete scores from 0,0.25,0.5,0.75,1{0,0.25,0.5,0.75,1}.
  • 0 (not natural / not similar)
  • 0.25 (mostly unnatural / mostly dissimilar)
  • 0.5 (neutral)
  • 0.75 (mostly natural / mostly similar)
  • 1 (fully natural / fully similar).
These annotations were performed by two independent annotators to ensure consistency and reliability. The aggregated results of this annotation study are summarized in Table 7.
We use Cohen’s Kappa to measure inter-annotator agreement across Naturalness (Nominal), Naturalness (Perturbed), and Semantic Similarity. Ratings were discretized into five ordinal bins following 5-point Likert scale (0, 0.25, 0.5, 0.75, 1.0) for categorical consistency. The kappa values show fair to moderate agreement (0.1-0.29), with slightly lower scores for similarity due to its subjective nature. Overall, these results indicate a consistent and reliable annotation process.
Table 7 presents the average annotation scores and inter-annotator agreement for Java, C++, and JavaScript. The results show that nominal prompts exhibit near-perfect naturalness (1.0), while perturbed prompts retain high naturalness (0.88–0.91), confirming that the perturbations remain largely fluent. Semantic similarity scores (0.94–0.96) indicate that meaning is well preserved across languages. The corresponding Cohen’s kappa values suggest fair to moderate agreement among annotators, reinforcing the consistency of the evaluation while reflecting the inherent subjectivity of semantic judgments in code understanding.
Table 7: Result of Human Annotations (Average Scores and Inter-Annotator Agreement) for 300 Random Samples
Scope Java C++ JavaScript
Naturalness (Nominal) $\uparrow$ 0.996 0.996 1.000
Naturalness (Perturbed) $\uparrow$ 0.914 0.888 0.900
Semantic Similarity $\uparrow$ 0.954 0.945 0.955
Cohen’s $\kappa$ (Nominal) 0.10 0.12 0.08
Cohen’s $\kappa$ (Perturbed) 0.22 0.29 0.15
Cohen’s $\kappa$ (Similarity) 0.26 0.18 0.06
@@ 5 Experiment Results
In this section, we aim to answer three research questions (RQs). For each RQ, we present the motivation and the approach to address the RQ, and the corresponding results.
Figure 6: Robust Drop among different languages and models
5.1 RQ1: How robust are LLMs across different programming languages?
@@
Motivation: Our experiment encompasses three distinct languages and four different models. A prior investigation (Wang et al., 2022) demonstrates the influence of perturbations on the robustness of code generation models. Nonetheless, the question of whether language-specific variations impact robustness remains unanswered. Hence, in this study, we aim to address this gap by exploring discrepancies in LLMs’ robustness across diverse programming languages (Java, C++, and JavaScript), shedding light on the extent of disparities among these languages.
@@
Approach: To investigate this research question, we conduct experiments on Java, C++, and JavaScript using 2 datasets: 1st is HumanEval-X, and 2nd is our curated dataset derived from HumanEval-X Zheng et al. (2023) and EvalPlus Liu et al. (2023). We evaluate six LLMs for code generation: Incoder-1B, Incoder-6B, CodeGen-Multi-2B, CodeGen-Multi-6B, Magicoder-7B, and QwenCode-3.5-6.7B. We evaluate the performance using three metrics: $RP_{s}@k$ measures robustness, $RD_{s}@k$ assesses the decrease in robustness due to changes, and $RR_{s}@k$ captures the relative differences between the original and modified datasets. The perturbations related to the Function Name and the DocString are compared with the nominal prompts (prompts that have only DocString and method signature), while the perturbations related to Syntax and Format are compared with the partial prompts (prompts that have some lines of code along with DocString and method signature).
@@
Results: Overall, the performance of models deteriorates after applying the perturbations. Table 8 and  9 present the worst-case performance of six models: Incoder-1B, Incoder-6B, CodeGen-2B-Multi, CodeGen-6B-Multi, Magicoder-7B, and QwenCode-3.5-6.7B under different perturbation types (DocString, Function Signature, Syntax, and Format) across three programming languages (Java, C++, and JavaScript). Similarly, Table 10 and  11 show the results in partial-case scenarios, where a model passes at least 3 out of 5 generated samples.
@@
Across all of the tables, we observe a consistent performance drop after applying perturbations. For instance, in Table 8, the metric $\text{RD}_{5}@1$, which quantifies the worst-case degradation, shows severe drops up to 0.9 (90%) in some DocString perturbation cases for Incoder-1B (Java). Even more superficial perturbations, such as Syntax and Format, lead to noticeable performance degradation, albeit to a lesser extent. Table 10 highlights that even under relaxed partial-pass criteria, the relative performance drop (RR5@1) remains substantial, typically ranging from 5% to 20%, depending on the model and language. This suggests that perturbations, whether semantic or syntactic, can significantly impact model behavior and reliability.
The performance drop of the LLMs is visualized in Figure 6. The figure shows the robust drop across four perturbation types: DocString, Syntax, Format, and Function Name, for all 6 models. It is organized into four columns (one per perturbation type) and six rows (one per model). Each plot represents a specific model–perturbation combination (e.g., Magicoder-7B with DocString) across three programming languages. Each plot contains three bars, one for each language. Higher bars indicate a larger robustness drop for that model under the corresponding perturbation.
@@
We observe a slight performance improvement in the partial-case scenario for Java under DocString and Function perturbations. This suggests that in some cases, modifying the function or the NL description can yield better results, hinting at the potential of prompt refinement as a strategy for improving code generation performance.
@@
In general, all models demonstrate performance degradation under perturbations, reaffirming the brittleness of current LLM-based code generation systems. While models such as Magicoder-7B show slightly better robustness, the consistent drop across all settings underscores the need for robustness-aware evaluation and training. Moreover, the variation in impact across programming languages suggests that the models’ sensitivity may be language-dependent, warranting further investigation into language-specific robustness strategies.
@@
If we see the language-wise trends, the three languages: Java, C++, and JavaScript exhibit distinct behaviors under input perturbations. Java shows the highest robustness, with notable RP5@1 gains and minimal RD5@1 drops, particularly under Format and Syntax perturbations. Its structured syntax likely helps models infer intent despite surface-level changes.
@@
Summary of Findings. The following trends can be observed from the results of RQ1.
  1. (1)